2. ​It is known that the weights of apples from a farm are normally distributed. In order to estimate the mean weight, a random sample of 150 apples is considered and the sample mean and population standard deviation are 6 kg and 0.8 kg respectively.

(a)​Construct a 95% confidence interval estimate for the population mean weight of apples.
​​​​​​​​​​​​ (2 marks)​​​​​​​​​​​​
(b)​The researcher suggests doing the study again so that 98% confidence interval estimate for the population mean weight of apples is (5.8835,6.1165) kg. How large should the sample size be?
​​​​​​​​​​​​ (3 marks)
1
Expert's answer
2020-12-01T03:04:22-0500

Given "\bar{x}=6\ kg, \sigma=0.8\ kg, n=150, \alpha=0.05"

(a)The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times \dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times \dfrac{\sigma}{\sqrt{n}})"

"=(6-1.96\times \dfrac{0.8}{\sqrt{150}}, 6+1.96\times \dfrac{0.8}{\sqrt{150}})"


"=(5.8720, 6.1280)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "5.8720<\mu<6.1280," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(5.8720, 6.1280)."


(b)The critical value for "\alpha=0.02" is "z_c=z_{1-\alpha/2}=2.326."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times \dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times \dfrac{\sigma}{\sqrt{n}})"

"=(5.8835, 6.1165)"

Then


"2\times z_c\times \dfrac{\sigma}{\sqrt{n}}=6.1165-5.8835"

"z_c\times \dfrac{\sigma}{\sqrt{n}}=0.1165"

"n=(z_c\times \dfrac{\sigma}{0.1165})^2"

"n=(2.326\times \dfrac{0.8}{0.1165})^2=255"

The sample size should be 255.